OpenStudy (anonymous):

For a large cooking project, Arthur needs to put 18 cups of water into a large pot. He can transport a maximum of 4 cups from the sink in a single trip. The pot already contains 5 cups of water. How many trips are required for Arthur to fill the pot to 18 cups? A. 3 B. 4 C. 5 D. 6

OpenStudy (anonymous):

\[ remaining = total - current \]

OpenStudy (anonymous):

\[total = 18 \]

OpenStudy (anonymous):

\[ current = 4 \]

OpenStudy (anonymous):

\[ remaining = 18-4 \]

OpenStudy (anonymous):

\[ \#\ trips = water \div water/trip \]

OpenStudy (anonymous):

\[ water/trip = 5 \]

OpenStudy (anonymous):

\[water = remaining = 18-4 \]

OpenStudy (anonymous):

\[ \#\ trips = (18-4)\div 5 \]